OpenStudy (anonymous):

Let f(x)=-2x+7 and g(x)=-6x+3. find f(x)=G(x) and state its domain possible answers are ...8x-4 ..-4x-10 ...8x-10 ...-4x-4

OpenStudy (anonymous):

f(6x+3) = -2(6x+3)+7 = -12x - 6 + 7 = -12x + 1

OpenStudy (anonymous):

-2x+7=-6x+3 -2x+4=-6x 4=-4x